### v1

@@ 9 Technical Appendix
    $\underline{k}_{i}=\underline{k}_{i}^{*}$ and $\bar{k}_{i}=\bar{k}_{i}^{*}$ if and only if $(\mathfrak{u}_{\underline{k}_{i}^{*}}-\mathfrak{u}_{k})\hat{p}\leq\tilde{\mathfrak{u}}_{k}-\tilde{\mathfrak{u}}_{\underline{k}_{i}^{*}}+({\mathfrak{u}}_{k}-{\mathfrak{u}}_{\underline{k}_{i}^{*}})\varepsilon_{i}$ for all $k=1,\ldots,K_{u}$ and $(\mathfrak{l}_{k}-\mathfrak{l}_{\bar{k}_{i}^{*}})\hat{p}\leq\tilde{\mathfrak{l}}_{\bar{k}_{i}^{*}}-\tilde{\mathfrak{l}}_{k}+({\mathfrak{l}}_{\bar{k}_{i}^{*}}-{\mathfrak{l}}_{k})\varepsilon_{i}$ for all $k=1,\ldots,K_{L}$;
@@
    $s_{i}^{\mathfrak{l}}=-$ and $s_{i}^{\mathfrak{u}}=-$ if and only if $\mathfrak{l}_{\bar{k}_{i}}\hat{p}\leq-\tilde{\mathfrak{l}}_{\bar{k}_{i}}-{\mathfrak{l}}_{\bar{k}_{i}}\varepsilon_{i}$ and $\mathfrak{u}_{\underline{k}_{i}}\hat{p}\leq-\tilde{\mathfrak{u}}_{\underline{k}_{i}}-{\mathfrak{u}}_{\underline{k}_{i}}\varepsilon_{i}$;
@@
    $\hat{d}=1$ if and only if $\sum_{i\in\underline{m}}(-{\mathfrak{u}}_{\underline{k}_{i}}\hat{p})+\sum_{i\in\bar{m}}(-{\mathfrak{l}}_{\bar{k}_{i}}\hat{p})\leq\sum_{i\in\underline{m}}(\tilde{\mathfrak{u}}_{\underline{k}_{i}}+{\mathfrak{u}}_{\underline{k}_{i}}\varepsilon_{i})+\sum_{i\in\bar{m}}(\tilde{\mathfrak{l}}_{\bar{k}_{i}}+{\mathfrak{l}}_{\bar{k}_{i}}\varepsilon_{i})$, where $\underline{m}=\{i\in\{1,\ldots,m\}:s_{i}^{\mathfrak{u}}=+\}$ and $\bar{m}=\{i\in\{1,\ldots,m\}:s_{i}^{\mathfrak{l}}=+\}.\quad\blacksquare$
Lemma 9.1.
Suppose Assumptions 3.13.5 hold. Then, for any $10<\alpha<$ 1,
@@
$\cdot\mathbb{P}\left(d\in\widehat{\mathcal{D}},\hat{j}_{L}(d)=j_{L}^{*},\hat{\gamma}_{L}(d)=\gamma_{L}^{*}\right)=0,$ (9.1)
@@
$\cdot\mathbb{P}\left(d\in\widehat{\mathcal{D}},\hat{j}_{U}(d)=j_{U}^{*},\hat{\gamma}_{U}(d)=\gamma_{U}^{*}\right)=0,$ (9.2)
for all $d\in\{0,\ldots,T\}$, $j_{L}^{*}\in\{1,\ldots,J_{L}\}$, $j_{U}^{*}\in\{1,\ldots,J_{U}\}$, $\gamma_{L}^{*}$ in the support of $\hat{\gamma}_{L}(d)$ and $\gamma_{U}^{*}$ in the support of $\hat{\gamma}_{U}(d)$.
Proof: The proof of (9.2) is nearly identical to the proof of (9.1) so that we only show the latter. Lemma A.1 of Lee et al., (2016) implies that $F_{TN}(t;\mu,\sigma^{2},v^{-},v^{+})$ is strictly decreasing in $\mu$ so that $\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p)\leq\widehat{L}(d)_{\alpha}^{C}$ is equivalent to
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p),\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d))\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d))\right)\right)\geq 1-\alpha,$
where we use $\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d))\right)$ and $\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d))\right)$ as shorthand for
$\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right)$
@@
$\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right).$
@@
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p),\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d))\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d))\right)\right)$
$=F_{TN}\left(\ell_{d,\hat{j}_{L}(d)}\sqrt{n}(\hat{p}-p);0,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}^{*}(d,\hat{j}_{L}(d))\right),\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}^{*}(d,\hat{j}_{L}(d))\right)\right),$
where $\widehat{Z}_{L}^{*}(d,j)=\sqrt{n}\hat{p}-\hat{b}_{L}(d,j)\ell_{d,j}\sqrt{n}(\hat{p}-p)$. Therefore, $\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p)\leq\widehat{L}(d)_{\alpha}^{C}$ is equivalent to
$F_{TN}\left(\ell_{d,\hat{j}_{L}(d)}\sqrt{n}(\hat{p}-p);0,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}^{*}(d,\hat{j}_{L}(d))\right),\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}^{*}(d,\hat{j}_{L}(d))\right)\right)\geq 1-\alpha.$ (9.3)
Under Assumptions 3.1, 3.3 and 3.5, a slight modification of Lemma 5 of Andrews et al., (2024) implies that to prove (9.1), it suffices to show that for all subsequences $\{n_{s}\}\subset\{n\}$, $\{\mathbb{P}_{n_{s}}\}\in\times_{n=1}^{\infty}\mathcal{P}_{n}$ with
@@
$\left(\sqrt{n_{s}}\ell_{d,j_{L}^{*}}(\hat{p}-p(\mathbb{P}_{n_{s}})),\widehat{\Sigma},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}^{*}(d,j_{L}^{*})\right),\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}^{*}(d,j_{L}^{*})\right)\right)$
@@
for all $d\in\{0,\ldots,T\}$, $j_{L}^{*}\in\{1,\ldots,J_{L}\}$ and $\gamma_{L}^{*}$ in the support of $\hat{\gamma}_{L}(d)$.
Using (9.4) and the equivalence in (9.3), the remaining arguments to prove (9.1) are nearly identical to those used in the proof of Proposition 1 of McCloskey, (2024) and therefore omitted for brevity. $\blacksquare$
Lemma 9.2.
Suppose Assumptions 3.13.5 hold. Then, for any $10<\alpha<$ 1,
@@
for all $d\in\{0,\ldots,T\}$.
Proof: The results of this lemma follow from Lemma 9.1 since, e.g.,
@@
Proof of Theorem 5.1: The result of this theorem follows from Lemma 9.2 since
@@
where the second inequality follows from the facts that $L(d)\geq\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p$ and $U(d)\leq\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}p$ almost surely and the final equality follows from Lemma 9.2. $\blacksquare$
@@
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\mathbb{P}\left(\sqrt{n}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)\geq\widehat{L}(\hat{d})_{\alpha}^{P}\right)\geq 1-\alpha.$ (9.5)
By the same argument as in the proof of Lemma 9.1, to prove (9.5), it suffices to show
$\lim_{n\rightarrow\infty}\mathbb{P}_{n_{s}}\left(\sqrt{n_{s}}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p(\mathbb{P}_{n_{s}}))\geq\widehat{L}(\hat{d})_{\alpha}^{P}\right)\geq 1-\alpha$ (9.6)
under conditions 1. and 3. Since $\sqrt{n_{s}}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)\geq\widehat{L}(\hat{d})_{\alpha}^{P}$ is equivalent to $\ell_{\hat{d},\hat{j}_{L}(\hat{d})}\sqrt{n_{s}}(\hat{p}-p)\leq\hat{c}_{1-\alpha,L}\sqrt{\widehat{\Sigma}_{L,\hat{d}J_{L}+\hat{j}_{L}(\hat{d})}}$, the left-hand side of (9.6) is equal to
@@
and $\Sigma_{L}=\ell^{mat}\Sigma\ell^{mat\prime}$, where all inequalities are taken element-wise across vectors, the inequality follows from the fact that $\ell_{\hat{d},\hat{j}_{L}(\hat{d})}\sqrt{n_{s}}(\hat{p}-p(\mathbb{P}_{n_{s}}))$ is a (random) element of $\ell^{mat}\sqrt{n_{s}}(\hat{p}-p(\mathbb{P}_{n_{s}}))$ and the first equality follows by identical arguments to those used in the proof of Proposition 11 of Andrews et al., (2024). We have thus proved (9.5). In addition,
@@
$\cdot\mathbb{P}\left(\hat{d}=d^{*},\sqrt{n}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)\geq\widehat{L}(\hat{d})_{\beta}^{P}\right)=0,$ (9.7)
@@
$\cdot\mathbb{P}(\hat{d}=d^{*},\sqrt{n}(\tilde{u}_{\hat{d},\hat{j}_{U}(\hat{d})}+u_{\hat{d},\hat{j}_{U}(\hat{d})}p)\leq\widehat{U}(\hat{d})_{1-\beta}^{P})=0,$ (9.8)
for all $d^{*}\in\{0,\ldots,T\}$.
Proof: The proof of (9.8) is nearly identical to the proof of (9.7) so that we only show the latter. Upon noting that $F_{TN}(t;\mu,\sigma^{2},\widehat{\mathcal{V}}_{L}^{-}(z,d,j,\gamma),\widehat{\mathcal{V}}_{L}^{+,H}(z,d,j,\gamma,\mu))$ is decreasing in $\mu$ by the same argument used in the proof of Proposition 5 of Andrews et al., (2024) and replacing condition 2. in the proof of Lemma 9.1 with
@@
completely analogous arguments to those used to prove (9.1) in Lemma 9.1 imply
@@
for all $d^{*}\in\{0,\ldots,T\}$, $j_{L}^{*}\in\{1,\ldots,J_{L}\}$ and $\gamma_{L}^{*}$ in the support of $\hat{\gamma}_{L}(d^{*})$. Then, the same argument as in the proof of Lemma 9.2 implies (9.7). $\blacksquare$
Lemma 9.4.
@@
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\mathbb{P}\left(\sqrt{n}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)>\widehat{L}(\hat{d})_{\alpha}^{H}\right)\geq 1-\alpha,$ (9.9)
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\mathbb{P}\left(\sqrt{n}(\tilde{u}_{\hat{d},\hat{j}_{U}(\hat{d})}+u_{\hat{d},\hat{j}_{U}(\hat{d})}p)<\widehat{U}(\hat{d})_{1-\alpha}^{H}\right)\geq 1-\alpha.$ (9.10)
Proof: The proof of (9.10) is nearly identical to the proof of (9.9) so that we only show the latter. Lemma 6 of Andrews et al., (2024) and Lemma 9.3 imply
@@
where the final inequality follows from (9.5) in the proof of Theorem 6.1. $\blacksquare$
Proof of Theorem 6.2: Using Lemma 9.4 in the place of Lemma 9.2, the proof is nearly identical to the proof of Theorem 5.1. $\blacksquare$

### v2

@@ Appendix C Technical Appendix
    $\underline{k}_{i}=\underline{k}_{i}^{*}$ and $\bar{k}_{i}=\bar{k}_{i}^{*}$ if and only if $(u_{\underline{k}_{i}^{*}}-u_{k})\hat{p}\leq\tilde{u}_{k}-\tilde{u}_{\underline{k}_{i}^{*}}+({u}_{k}-{u}_{\underline{k}_{i}^{*}})\varepsilon_{i}$ for all $k=1,\ldots,J_{u}$ and $(\ell_{k}-\ell_{\bar{k}_{i}^{*}})\hat{p}\leq\tilde{\ell}_{\bar{k}_{i}^{*}}-\tilde{\ell}_{k}+({\ell}_{\bar{k}_{i}^{*}}-{\ell}_{k})\varepsilon_{i}$ for all $k=1,\ldots,J_{L}$;
@@
    $s_{i}^{\ell}=-$ and $s_{i}^{u}=-$ if and only if $\ell_{\bar{k}_{i}}\hat{p}\leq-\tilde{\ell}_{\bar{k}_{i}}-{\ell}_{\bar{k}_{i}}\varepsilon_{i}$ and $u_{\underline{k}_{i}}\hat{p}\leq-\tilde{u}_{\underline{k}_{i}}-{u}_{\underline{k}_{i}}\varepsilon_{i}$;
@@
    $\hat{d}=1$ if and only if $\sum_{i\in\underline{m}}(-{u}_{\underline{k}_{i}}\hat{p})+\sum_{i\in\bar{m}}(-{\ell}_{\bar{k}_{i}}\hat{p})\leq\sum_{i\in\underline{m}}(\tilde{u}_{\underline{k}_{i}}+{u}_{\underline{k}_{i}}\varepsilon_{i})+\sum_{i\in\bar{m}}(\tilde{\ell}_{\bar{k}_{i}}+{\ell}_{\bar{k}_{i}}\varepsilon_{i})$, where $\underline{m}=\{i\in\{1,\ldots,m\}:s_{i}^{u}=+\}$ and $\bar{m}=\{i\in\{1,\ldots,m\}:s_{i}^{\ell}=+\}.\quad\blacksquare$
Lemma C.1.
Suppose Assumptions 3.1, 3.2 and 3.43.6 hold. Then, for any $10<\alpha<$ 1,
@@
$\cdot\mathbb{P}\left(d\in\widehat{\mathcal{D}},\hat{j}_{L}(d)=j_{L}^{*},\hat{\gamma}_{L}(d)=\gamma_{L}^{*}\right)=0,$ (C.1)
@@
$\cdot\mathbb{P}\left(d\in\widehat{\mathcal{D}},\hat{j}_{U}(d)=j_{U}^{*},\hat{\gamma}_{U}(d)=\gamma_{U}^{*}\right)=0,$ (C.2)
for all $d\in\{d^{0},\ldots,d^{K}\}$, $j_{L}^{*}\in\{1,\ldots,J_{L}\}$, $j_{U}^{*}\in\{1,\ldots,J_{U}\}$, $\gamma_{L}^{*}$ in the support of $\hat{\gamma}_{L}(d)$ and $\gamma_{U}^{*}$ in the support of $\hat{\gamma}_{U}(d)$.
Proof: The proof of (C.2) is nearly identical to the proof of (C.1) so that we only show the latter. Lemma A.1 of Lee et al., (2016) implies that $F_{TN}(t;\mu,\sigma^{2},v^{-},v^{+})$ is strictly decreasing in $\mu$ so that $\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p)\leq\widehat{L}(d)_{\alpha}^{C}$ is equivalent to
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p),\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime}\right|\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d))\right),$
$\left.\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d))\right)\right)\geq 1-\alpha,$
where we use $\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d))\right)$ and $\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d))\right)$ as shorthand for
$\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right)$
@@
$\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right).$
@@
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p),\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime}\right|\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d))\right),$
$\left.\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d))\right)\right)$
$=F_{TN}\left(\ell_{d,\hat{j}_{L}(d)}\sqrt{n}(\hat{p}-p);0,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime}\right|\left.\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}^{*}(d,\hat{j}_{L}(d))\right),\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}^{*}(d,\hat{j}_{L}(d))\right)\right),$
where $\widehat{\mathcal{Z}}_{L}^{*}(d,j)=\sqrt{n}\hat{p}-\hat{b}_{L}(d,j)\ell_{d,j}\sqrt{n}(\hat{p}-p)$. Therefore, $\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p)\leq\widehat{L}(d)_{\alpha}^{C}$ is equivalent to
$F_{TN}\left(\ell_{d,\hat{j}_{L}(d)}\sqrt{n}(\hat{p}-p);0,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime}\right|\left.\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}^{*}(d,\hat{j}_{L}(d))\right),\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}^{*}(d,\hat{j}_{L}(d))\right)\right)\geq 1-\alpha.$ (C.3)
Under Assumptions 3.1, 3.4 and 3.6, a slight modification of Lemma 5 of Andrews et al., (2024) implies that to prove (C.1), it suffices to show that for all subsequences $\{n_{s}\}\subset\{n\}$, $\{\mathbb{P}_{n_{s}}\}\in\times_{n=1}^{\infty}\mathcal{P}_{n}$ with
@@
$\left(\sqrt{n_{s}}\ell_{d,j_{L}^{*}}(\hat{p}-p(\mathbb{P}_{n_{s}})),\widehat{\Sigma},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}^{*}(d,j_{L}^{*})\right),\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}^{*}(d,j_{L}^{*})\right)\right)$
@@
for all $d\in\{d^{0},\ldots,d^{K}\}$, $j_{L}^{*}\in\{1,\ldots,J_{L}\}$ and $\gamma_{L}^{*}$ in the support of $\hat{\gamma}_{L}(d)$.
Using (C.4) and the equivalence in (C.3), the remaining arguments to prove (C.1) are nearly identical to those used in the proof of Proposition 1 of McCloskey, (2024) and therefore omitted for brevity. $\blacksquare$
Lemma C.2.
Suppose Assumptions 3.1, 3.2 and 3.43.6 hold. Then, for any $10<\alpha<$ 1,
@@
for all $d\in\{d^{0},\ldots,d^{K}\}$.
Proof: The results of this lemma follow from Lemma C.1 since, e.g.,
@@
Proof of Theorem 5.1: The result of this theorem follows from Lemma C.2 since
@@
where the second inequality follows from the facts that $L(d)\geq\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}p$ and $U(d)\leq\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}p$ almost surely and the final equality follows from Lemma C.2. $\blacksquare$
@@
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\mathbb{P}\left(\sqrt{n}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)\geq\widehat{L}(\hat{d})_{\alpha}^{P}\right)\geq 1-\alpha.$ (C.5)
By the same argument as in the proof of Lemma C.1, to prove (C.5), it suffices to show
$\lim_{n\rightarrow\infty}\mathbb{P}_{n_{s}}\left(\sqrt{n_{s}}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p(\mathbb{P}_{n_{s}}))\geq\widehat{L}(\hat{d})_{\alpha}^{P}\right)\geq 1-\alpha$ (C.6)
under conditions 1. and 3. Since $\sqrt{n_{s}}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)\geq\widehat{L}(\hat{d})_{\alpha}^{P}$ is equivalent to $\ell_{\hat{d},\hat{j}_{L}(\hat{d})}\sqrt{n_{s}}(\hat{p}-p)\leq\hat{c}_{1-\alpha,L}\sqrt{\widehat{\Sigma}_{L,\hat{d}J_{L}+\hat{j}_{L}(\hat{d})}}$, the left-hand side of (C.6) is equal to
@@
and $\Sigma_{L}=\ell^{mat}\Sigma\ell^{mat\prime}$, where all inequalities are taken element-wise across vectors, the inequality follows from the fact that $\ell_{\hat{d},\hat{j}_{L}(\hat{d})}\sqrt{n_{s}}(\hat{p}-p(\mathbb{P}_{n_{s}}))$ is a (random) element of $\ell^{mat}\sqrt{n_{s}}(\hat{p}-p(\mathbb{P}_{n_{s}}))$ and the first equality follows by identical arguments to those used in the proof of Proposition 11 of Andrews et al., (2024). We have thus proved (C.5). In addition,
@@
$\cdot\mathbb{P}\left(\hat{d}=d^{*},\sqrt{n}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)\geq\widehat{L}(\hat{d})_{\beta}^{P}\right)=0,$ (C.7)
@@
$\cdot\mathbb{P}(\hat{d}=d^{*},\sqrt{n}(\tilde{u}_{\hat{d},\hat{j}_{U}(\hat{d})}+u_{\hat{d},\hat{j}_{U}(\hat{d})}p)\leq\widehat{U}(\hat{d})_{1-\beta}^{P})=0,$ (C.8)
for all $d^{*}\in\{d^{0},\ldots,d^{K}\}$.
Proof: The proof of (C.8) is nearly identical to the proof of (C.7) so that we only show the latter. Upon noting that $F_{TN}(t;\mu,\sigma^{2},\widehat{\mathcal{V}}_{L}^{-}(z,d,j,\gamma),\widehat{\mathcal{V}}_{L}^{+,H}(z,d,j,\gamma,\mu))$ is decreasing in $\mu$ by the same argument used in the proof of Proposition 5 of Andrews et al., (2024) and replacing condition 2. in the proof of Lemma C.1 with
@@
completely analogous arguments to those used to prove (C.1) in Lemma C.1 imply
@@
for all $d^{*}\in\{d^{0},\ldots,d^{K}\}$, $j_{L}^{*}\in\{1,\ldots,J_{L}\}$ and $\gamma_{L}^{*}$ in the support of $\hat{\gamma}_{L}(d^{*})$. Then, the same argument as in the proof of Lemma C.2 implies (C.7). $\blacksquare$
Lemma C.4.
@@
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\mathbb{P}\left(\sqrt{n}(\tilde{\ell}_{\hat{d},\hat{j}_{L}(\hat{d})}+\ell_{\hat{d},\hat{j}_{L}(\hat{d})}p)>\widehat{L}(\hat{d})_{\alpha}^{H}\right)\geq 1-\alpha,$ (C.9)
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\mathbb{P}\left(\sqrt{n}(\tilde{u}_{\hat{d},\hat{j}_{U}(\hat{d})}+u_{\hat{d},\hat{j}_{U}(\hat{d})}p)<\widehat{U}(\hat{d})_{1-\alpha}^{H}\right)\geq 1-\alpha.$ (C.10)
Proof: The proof of (C.10) is nearly identical to the proof of (C.9) so that we only show the latter. Lemma 6 of Andrews et al., (2024) and Lemma C.3 imply
@@
where the final inequality follows from (C.5) in the proof of Theorem 5.2. $\blacksquare$
Proof of Theorem 5.3: Using Lemma C.4 in the place of Lemma C.2, the proof is nearly identical to the proof of Theorem 5.1. $\blacksquare$